OpenStudy (anonymous):

the polynomial x^3+5x^2-57x-189 expresses the volume in cubic inches of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions. (hint: the height is greater than the depth)

OpenStudy (anonymous):

first since given x+3 =15 then x=12 now we have to find other factors of x^3+5x^2-57x-189 as one of the factor is x+3 so u can divide to get the other facots now x^3+5x^2-57x-189 =x^3+3x^2 +2x^2 +6x -63x -189 =x^2(x+3) +2x(x+3) -63(x+3) =(x+3)(x^2+2x-63) =(x+3)(x^2+9x-7x-63) =(x+3)(x(x+9)-7(x+9)) =(x+3)(x-7)(x+9)

OpenStudy (anonymous):

thus the other sides might be one of these (x+9) and (x-7) since the height is greater than the depth height is 12+9 =21 and the length is 12-7=5

OpenStudy (anonymous):

omg thanks so much

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

but is it clear to understand